# Masked Language Modeling Becomes Conditional Density Estimation for Tabular Data Synthesis

### v1

@@ 3.3 Results
We attribute the effectiveness of the feature selection downstream task to our emphasis on estimating ‘conditional’ distributions. In Random Forest [4], each node in a decision tree represents a conditional distribution of a variable conditioned on the splits made by the tree up to that node, and the feature importance is determined by the purity of nodes. Therefore, accurately estimating the conditional distribution can lead to higher performance in preserving the feature importance ranking.
Q2. The missing data mechanism within the parentheses refers to the mechanism applied to the training dataset on which MaCoDE was trained. Despite encountering missing data scenarios such as MAR, MCAR, MNARL, and MNARQ, Table 3.3 illustrates MaCoDE’s ability to generate high-quality synthetic data regarding machine learning utility while handling incomplete training datasets without significant performance degradation. Even in the presence of missing entries, MaCoDE achieves better metric scores than most baseline models in terms of SMAPE, except for TabDDPM. Additionally, concerning the $F_{1}$ score, MaCoDE either competes competitively or achieves a higher score than other baseline models.
Table 2: Q2: Machine learning utility. The means and standard errors of the mean across 10 datasets and 10 repeated experiments are reported. ‘Baseline’ refers to the result obtained using half of the real training dataset. Across all missingness patterns, a missingness rate of 0.3 is employed. $\uparrow$ ($\downarrow$) denotes higher (lower) is better. The best value is bolded, and the second best is underlined.
Model SMAPE $\downarrow$ $F_{1}$uparrow
Baseline .$107_{\pm.008}.$$686_{\pm.023}.$
CTGAN .$256_{\pm.016}.$$411_{\pm.027}.$
TVAE .$192_{\pm.011}.$$608_{\pm.021}.$
CTAB-GAN .$218_{\pm.012}.$$524_{\pm.026}.$
CTAB-GAN+ .$226_{\pm.017}.$$530_{\pm.020}.$
DistVAE .$226_{\pm.017}.$$588_{\pm.021}.$
TabDDPM .$\underline{.161}_{\pm.011}$$576_{\pm.022}.$
TabMT .$188_{\pm.013}.$$622_{\pm.024}.$
MaCoDE .$\mathbf{.158}_{\pm.010}$$\mathbf{.635}_{\pm.023}$
MaCoDE(MCAR) .$168_{\pm.011}.$$623_{\pm.023}.$
MaCoDE(MAR) .$167_{\pm.011}.$$626_{\pm.023}.$
MaCoDE(MNARL) .$169_{\pm.011}.$$624_{\pm.023}.$
MaCoDE(MNARQ) .$164_{\pm.010}.$$\underline{.630}_{\pm.023}$
Table 3: Q3: Multiple imputation under MAR at 0.3 missingness. The means and standard errors of the mean across 10 datasets and 10 repeated experiments are reported. $\downarrow$ denotes lower is better. Coverage close to 0.95 indicates better performance. The best value is bolded, and the second best is underlined.
@@
MICE .$011_{\pm.001}.$$750_{\pm.023}.$$\textbf{.036}_{\pm.002}.$
GAIN .$026_{\pm.002}.$$574_{\pm.024}.$$\textbf{.036}_{\pm.002}.$
missMDA1 .$017_{\pm.001}.$$705_{\pm.025}.$$047_{\pm.002}.$
VAEAC .$\underline{.008}_{\pm.000}$$785_{\pm.022}.$$\textbf{.036}_{\pm.002}.$
MIWAE .$\textbf{.007}_{\pm.000}.$$\underline{.865}_{\pm.018}$$040_{\pm.002}.$
not-MIWAE .$\underline{.008}_{\pm.001}$$844_{\pm.020}.$$\underline{.038}_{\pm.002}$
EGC2 .$\textbf{.007}_{\pm.000}.$$\textbf{.980}_{\pm.003}.$$064_{\pm.001}.$
MaCoDE(MAR) .$010_{\pm.001}.$$826_{\pm.023}.$$\underline{.038}_{\pm.002}$
  • 1We faced out-of-memory issues with covtype and letter datasets using the missMDA package in R. We excluded these datasets from the results. See Appendix A.7.5 for dataset-specific results.
  • 2We found that concrete, kings, and loan datasets are not compatible with gcimpute package. We excluded these datasets from the results. See Appendix A.7.6 for dataset-specific results.
@@
Q3. The missing data mechanism within the parentheses refers to the mechanism applied to the dataset on which MaCoDE was trained. Table 3.3 indicates that MaCoDE consistently exhibits competitive performance against all baseline models across metrics assessing multiple imputation performances, including bias, coverage, and confidence interval length. This suggests our proposed approach can support multiple imputations for deriving statistically valid inferences from missing data with the MAR mechanism.
(a) Q2 under MAR.
(b) Q3 under MAR.
@@
Figure 3: Sensitivity analysis with respect to missingness rate using kings dataset. (a): machine learning utility. (b): multiple imputation. The means and the standard errors of the mean across 10 repeated experiments are reported. Error bars represent standard errors.
@@
Sensitivity analysis. We also conducted a sensitivity analysis by varying the missingness rate of kings dataset. Figure 3.(a) illustrates that, in terms of SMAPE, MaCoDE maintains competitive performance even as the missingness rate increases. Concerning the $F_{1}$ score, MaCoDE outperforms other models at missingness rates of 0.1 and 0.3, but its performance declines beyond a missingness rate of 0.5. Additionally, regarding multiple imputation performance of Figure 3.(b), despite the increase in missingness rate, MaCoDE consistently demonstrates competitive performance compared to other imputation models without significant performance degradation.
4 Related Works
Synthetic tabular data generation. Deep generative models such as Variational Autoencoders (VAEs) and Generative Adversarial Networks (GANs) have been employed in synthetic tabular data generation. CTGAN [51], TVAE [51], CTAB-GAN [56], CTAB-GAN+ [57], DistVAE [1], and GOGGLE [31] employ VAE or GAN for their ability to represent complex data distributions. Recently, diffusion-based or score-based generative models, such as TabDDPM [24], STaSy [23], and CoDi [25], have gained popularity due to their noticeable performance in approximating data distributions. However, these models often encounter challenges when dealing with incomplete training datasets as they rely on complete data for training. In contrast, Transformer-based synthesizers like TabPFGen [33], TabMT [13], and REaLTabFormer [45] can handle incomplete datasets.
Missing data imputations. Recent methods employ the deep generative model to estimate a joint distribution and generate samples for imputations. GAN-based imputers such as GAIN [52] and MisGAN [30] adopt an adversarial learning approach to generate both missing entries and masking vectors. Other approaches like VAEAC [19], HI-VAE [37], and ReMasker [10] employ strategies that allow them to learn conditional distributions on arbitrary conditioning sets using the uniform masking strategy. MIWAE [34], based on the Importance Weighted Autoencoder [5], demonstrates that under mild conditions and the MAR assumption, the target likelihood can be approximated regardless of the imputation function. Additionally, not-MIWAE [18] extends MIWAE to handle cases where the data is under MNAR assumption by modeling missing entries as a latent variable. In parallel, the Optimal Transport (OT)-based method employs distributional matching, utilizing the 2-Wasserstein distance to compare distributions in both data and latent spaces [36, 54]. To handle mixed-type tabular datasets, [55] introduced EGC (extended Gaussian copula), which relies on a latent Gaussian distribution to support single and multiple imputations.
@@ 5 Conclusions and Limitations
This paper introduces a novel approach to generating synthetic data for mixed-type tabular datasets. Our proposed method integrates histogram-based non-parametric conditional density estimation and the MLM-based approach. By demonstrating that our conditional density estimators for continuous columns are weakly consistent estimators, we bridge the theoretical gap between distributional learning and the multi-class classification task of MLM. Although our primary goal is to generate synthetic data with high MLu, we empirically demonstrate that we achieve high joint statistical fidelity and MLu simultaneously. Furthermore, empirical experiments validate that our proposed model can generate high-quality synthetic tabular datasets in terms of MLu even when incomplete training datasets are given.
However, in our context, ‘arbitrary’ conditional density estimation refers to estimating the conditional density for arbitrary combinations of conditioning sets and target variables rather than accommodating arbitrary types of continuous distributions. For example, if ${\bf x}$ has lower bounded support and its CDF doesn’t satisfy Assumption 1, then $Q(0)=\inf\{{\bf x}\in\mathbb{R}:0\leq F({\bf x})\}=-\infty$, indicating that inverse transform sampling may generate out-of-support samples. Improving our proposed model to accommodate arbitrary types of continuous distributions is our future work.
@@ A.1 Proof of Theorem 1
This proof is based on Theorem 6.11 of [50] and Theorem 1 of [29].
Without loss of generality, let $\sigma$ be an identity permutation. We equally space grid points on the interval $[0,1]$ such that bins are represented as $[b_{l-1},b_{l}),l=1,\cdots,L_{n}$, where $h$ denotes the bin width such that $L_{n}h=1$. We consider arbitrary $j\in I_{C}$ and ${\bf x}_{k}\in{\cal X}_{k},k=1,\cdots,j$ such that $p^{*}_{j}({\bf x}_{j})>0$. Note that ${\bf x}_{1},{\bf x}_{2},\cdots,{\bf x}_{j}$ are given and fixed. For notational simplicity, let ${\bf x}_{-j}\coloneqq({\bf x}_{1},\cdots,{\bf x}_{j-1})$.
We partition the $n$ training samples into two equal halves, each containing $n/2$ training samples, and denote the data distributions derived from the two equal halves $\{{\bf x}^{(i)}\}_{i=1}^{n/2}$ and $\{{\bf x}^{(i)}\}_{i=n/2+1}^{n}$ as $\mathbb{P}_{1}$ and $\mathbb{P}_{2}$, respectively. The expectations, variances, and biases subscripted by $\mathbb{P}_{1}$ and $\mathbb{P}_{2}$ indicate that they depend on these respective data distributions. If no subscript is present, it implies they range over the data distribution derived from the entire dataset, $\{{\bf x}^{(i)}\}_{i=1}^{n}$.
The ground-truth conditional probability of $g({\bf x};F^{*})_{j}=l$ given ${\bf x}_{-j}$ is defined as
$\Pr\left(g({\bf x};F^{*})_{j}=l\big{|}{\bf x}_{-j}\right)\eqqcolon\pi^{*}_{jl}\big{(}{\bf x}_{-j}\big{)},$
which is approximated by the model, $\pi_{jl}(g({\bf x};\hat{F})\odot{\bf m}^{(j)};\theta)$. It’s crucial to note that $\hat{F}$ and the neural network parameters $\theta$ (i.e., $\pi_{j}(\cdot;\theta)$) are determined through training and fitting on the first half of the dataset, implying their dependency on $\mathbb{P}_{1}$.
We define the bias and variance of $\pi_{jl}$ as
$\mbox{Bias}\Big{(}\pi_{jl}(g({\bf x};\hat{F})\odot{\bf m}^{(j)};\theta)\Big{)}$$\coloneqq$ $\mathbb{E}_{\mathbb{P}_{1}}\Big{[}\pi_{jl}(g({\bf x};\hat{F})\odot{\bf m}^{(j)};\theta)\Big{]}-\pi^{*}_{jl}({\bf x}_{-j})$
$\mbox{Var}\Big{(}\pi_{jl}(g({\bf x};\hat{F})\odot{\bf m}^{(j)};\theta)\Big{)}$$\coloneqq$ $\mathbb{E}_{\mathbb{P}_{1}}\Bigg{[}\Big{(}\pi_{jl}(g({\bf x};\hat{F})\odot{\bf m}^{(j)};\theta)-\mathbb{E}_{\mathbb{P}_{1}}[\pi_{jl}(g({\bf x};\hat{F})\odot{\bf m}^{(j)};\theta)]\Big{)}^{2}\Bigg{]},$
which are denoted as $\mbox{Bias}(\pi_{jl})$ and $\mbox{Var}(\pi_{jl})$, respectively, for notational simplicity.
We have
$\mathbb{E}\Big{[}\hat{c}_{j}(\tilde{F}_{j}({\bf x}_{j})|{\bf x}_{-j};\theta)\Big{]}$ $=$ $\mathbb{E}\left[\sum_{l=1}^{L_{n}}\frac{\mathbb{I}(\tilde{F}_{j}({\bf x}_{j})\in[b_{l-1},b_{l}))}{1/{L_{n}}}\cdot\pi_{jl}(g({\bf x};\hat{F})\odot{\bf m}^{(j)};\theta)\right]$ (7)
$=$ $\sum_{l=1}^{L_{n}}\mathbb{E}_{\mathbb{P}_{2}}[\mathbb{I}(\tilde{F}_{j}({\bf x}_{j})\in[b_{l-1},b_{l}))]\cdot\frac{\mathbb{E}_{\mathbb{P}_{1}}[\pi_{jl}(g({\bf x};\hat{F})\odot{\bf m}^{(j)};\theta)]}{h}$
$=$ $\sum_{l=1}^{L_{n}}\mathbb{E}_{\mathbb{P}_{2}}[\mathbb{I}(\tilde{F}_{j}({\bf x}_{j})\in[b_{l-1},b_{l}))]\cdot\left(\frac{\pi_{jl}^{*}({\bf x}_{-j})+\mbox{Bias}(\pi_{jl})}{h}\right)$
where the first equality follows from the definition of $\hat{c}_{j}$, and the second equality follows from the independence between ($\pi_{j}(\cdot;\theta)$, $\hat{F}_{j}$) and $\tilde{F}_{j}$ under Assumption 3.
Under Assumption 2, $\pi_{jl}^{*}({\bf x}_{-j})$ is written as:
$\pi_{jl}^{*}({\bf x}_{-j})$
@@
$=$ $h\cdot c^{*}_{j}(F^{*}_{j}({\bf x}_{j})|{\bf x}_{-j})+h\cdot{{c_{j}^{*}}^{\prime}}(F^{*}_{j}({\bf x}_{j})|{\bf x}_{-j})(v_{l}^{(0)}-F^{*}_{j}({\bf x}_{j}))+\mathcal{O}(h^{3})$
by Taylor’s theorem, for some $u_{j}^{*}$ between $v$ and $F^{*}_{j}({\bf x}_{j})$, where $v_{l}^{(0)}$ is the middle point of interval $[b_{l-1},b_{l})$.
Thus, by (7), the bias for the conditional density estimator $\hat{c}_{j}$ is written as
$\mbox{Bias}\Big{(}\hat{c}_{j}(\tilde{F}_{j}({\bf x}_{j})|{\bf x}_{-j};\theta)\Big{)}$
$=$ $\mathbb{E}\Big{[}\hat{c}_{j}(\tilde{F}_{j}({\bf x}_{j})|{\bf x}_{-j};\theta)\Big{]}-c^{*}_{j}(F^{*}_{j}({\bf x}_{j})|{\bf x}_{-j})$
$=$ $\sum_{l=1}^{L_{n}}\mathbb{E}_{\mathbb{P}_{2}}[\mathbb{I}(\tilde{F}_{j}({\bf x}_{j})\in[b_{l-1},b_{l}))]\cdot\left(\frac{\pi_{jl}^{*}({\bf x}_{-j})+\mbox{Bias}(\pi_{jl})}{h}\right)-c^{*}_{j}(F^{*}_{j}({\bf x}_{j})|{\bf x}_{-j})$
$=$ $\sum_{l=1}^{L_{n}}\mathbb{E}_{\mathbb{P}_{2}}[\mathbb{I}(\tilde{F}_{j}({\bf x}_{j})\in[b_{l-1},b_{l}))]\cdot\Bigg{(}\frac{\mbox{Bias}(\pi_{jl})}{h}+{{c_{j}^{*}}^{\prime}}(F^{*}_{j}({\bf x}_{j})|{\bf x}_{-j})(v_{l}^{(0)}-F^{*}_{j}({\bf x}_{j}))+\mathcal{O}(h^{2})\Bigg{)}$
$\leq$ $\sup_{l\in[L_{n}]}\Bigg{(}\frac{\mbox{Bias}(\pi_{jl})}{h}+{{c_{j}^{*}}^{\prime}}(F^{*}_{j}({\bf x}_{j})|{\bf x}_{-j})(v_{l}^{(0)}-F^{*}_{j}({\bf x}_{j}))+\mathcal{O}(h^{2})\Bigg{)}$
since $\sum_{l=1}^{L_{n}}\mathbb{E}_{\mathbb{P}_{2}}[\mathbb{I}(\tilde{F}_{j}({\bf x}_{j})\in[b_{l-1},b_{l}))]=1$. And
$\mbox{Bias}^{2}(\hat{c}_{j})\leq 2\cdot\sup_{l\in[L_{n}]}\Bigg{(}\frac{\mbox{Bias}^{2}(\pi_{jl})}{h^{2}}+({{c_{j}^{*}}^{\prime}}(F^{*}_{j}({\bf x}_{j})|{\bf x}_{-j}))^{2}(v_{l}^{(0)}-F^{*}_{j}({\bf x}_{j}))^{2}+\mathcal{O}(h^{4})\Bigg{)},$ (8)
where $\mbox{Bias}(\hat{c}_{j})$ denotes $\mbox{Bias}\big{(}\hat{c}_{j}({\bf u}_{j}|{\bf x}_{-j};\theta)\big{)}$.
Since $\mathbb{I}(\tilde{F}_{j}({\bf x}_{j})\in[b_{l-1},b_{l}))\cdot\mathbb{I}(\tilde{F}_{j}({\bf x}_{j})\in[b_{k-1},b_{k}))=0$ for $l\neq k$, and $\mathbb{E}_{\mathbb{P}_{2}}[\mathbb{I}(\tilde{F}_{j}({\bf x}_{j})\in[b_{l-1},b_{l}))]\leq 1$, and $\mathbb{E}_{\mathbb{P}_{1}}[\pi_{jl}(g({\bf x};\hat{F})\odot{\bf m}^{(j)};\theta)]\leq 1$, the variance for the conditional density estimator $\hat{c}_{j}$ is written as
$\mbox{Var}\Big{(}\hat{c}_{j}(\tilde{F}_{j}({\bf x}_{j})|{\bf x}_{-j};\theta)\Big{)}$
$=$ $\mbox{Var}\left(\sum_{l=1}^{L_{n}}\frac{\mathbb{I}(\tilde{F}_{j}({\bf x}_{j})\in[b_{l-1},b_{l}))}{1/L_{n}}\cdot\pi_{jl}(g({\bf x};\hat{F})\odot{\bf m}^{(j)};\theta)\right)$
$\leq$ $\frac{1}{h^{2}}\left(\sum_{l=1}^{L_{n}}\mbox{Var}\left(\mathbb{I}(\tilde{F}_{j}({\bf x}_{j})\in[b_{l-1},b_{l}))\cdot\pi_{jl}(g({\bf x};\hat{F})\odot{\bf m}^{(j)};\theta)\right)\right)$
$\leq$ $\frac{1}{h^{2}}\left(L_{n}\cdot\sup_{l\in[L_{n}]}\mbox{Var}\left(\mathbb{I}(\tilde{F}_{j}({\bf x}_{j})\in[b_{l-1},b_{l}))\cdot\pi_{jl}(g({\bf x};\hat{F})\odot{\bf m}^{(j)};\theta)\right)\right)$
$=$ $\frac{1}{h^{3}}\cdot\sup_{l\in[L_{n}]}\Bigg{\{}\mbox{Var}\Big{(}\mathbb{I}(\tilde{F}_{j}({\bf x}_{j})\in[b_{l-1},b_{l}))\Big{)}\cdot\mbox{Var}\Big{(}\pi_{jl}(g({\bf x};\hat{F})\odot{\bf m}^{(j)};\theta)\Big{)}$
$+\mbox{Var}\Big{(}\mathbb{I}(\tilde{F}_{j}({\bf x}_{j})\in[b_{l-1},b_{l}))\Big{)}\cdot\left(\mathbb{E}_{\mathbb{P}_{1}}\Big{[}\pi_{jl}(g({\bf x};\hat{F})\odot{\bf m}^{(j)};\theta)\Big{]}\right)^{2}$
$+\left(\mathbb{E}_{\mathbb{P}_{2}}\Big{[}\mathbb{I}(\tilde{F}_{j}({\bf x}_{j})\in[b_{l-1},b_{l}))\Big{]}\right)^{2}\cdot\mbox{Var}\Big{(}\pi_{jl}(g({\bf x};\hat{F})\odot{\bf m}^{(j)};\theta)\Big{)}\Bigg{\}}$
$\leq$ $\frac{1}{h^{3}}\cdot\sup_{l\in[L_{n}]}\Bigg{\{}\mbox{Var}(\pi_{jl})\cdot\mbox{Var}\Big{(}\mathbb{I}(\tilde{F}_{j}({\bf x}_{j})\in[b_{l-1},b_{l}))\Big{)}+\mbox{Var}(\pi_{jl})+\mbox{Var}\Big{(}\mathbb{I}(\tilde{F}_{j}({\bf x}_{j})\in[b_{l-1},b_{l}))\Big{)}\Bigg{\}},$
where the first inequality follows from the Cauchy-Schwarz inequality and the second equality follows from the independence between ($\pi_{j}(\cdot;\theta)$, $\hat{F}_{j}$) and $\tilde{F}_{j}$ under Assumption 3. For notational simplicity, we denote $\mbox{Var}\Big{(}\hat{c}_{j}(\tilde{F}_{j}({\bf x}_{j})|{\bf x}_{-j};\theta)\Big{)}$ as $\mbox{Var}(\hat{c}_{j})$.
Therefore, by (8) and (A.1), the MSE (mean squared error) of the conditional density estimator $\hat{c}_{j}$ is written as follows:
$\mbox{MSE}\Big{(}\hat{c}_{j}(\tilde{F}_{j}({\bf x}_{j})|{\bf x}_{-j};\theta)\Big{)}=\mbox{Bias}^{2}(\hat{c}_{j})+\mbox{Var}(\hat{c}_{j})$
$\leq$ $2\cdot\sup_{l\in[L_{n}]}\Biggl{(}\frac{\mbox{Bias}^{2}(\pi_{jl})}{h^{2}}+({{c_{j}^{*}}^{\prime}}(F^{*}_{j}({\bf x}_{j})|{\bf x}_{-j}))^{2}(v_{l}^{(0)}-F^{*}_{j}({\bf x}_{j}))^{2}+\mathcal{O}(h^{3})\Biggl{)}$
$+\sup_{l\in[L_{n}]}\Biggl{(}\frac{\mbox{Var}(\pi_{jl})\cdot\mbox{Var}\Big{(}\mathbb{I}(\tilde{F}_{j}({\bf x}_{j})\in[b_{l-1},b_{l}))\Big{)}+\mbox{Var}(\pi_{jl})+\mbox{Var}\Big{(}\mathbb{I}(\tilde{F}_{j}({\bf x}_{j})\in[b_{l-1},b_{l}))\Big{)}}{h^{3}}\Biggl{)}$
$\leq$ $2\cdot\frac{\Big{(}\sup_{l\in[L_{n}]}\mbox{Bias}(\pi_{jl})\Big{)}^{2}}{h^{2}}+\frac{\sup_{l\in[L_{n}]}\mbox{Var}(\pi_{jl})\cdot\mbox{Var}\Big{(}\mathbb{I}(\tilde{F}_{j}({\bf x}_{j})\in[b_{l-1},b_{l}))\Big{)}}{h^{3}}$
$+\frac{\sup_{l\in[L_{n}]}\mbox{Var}(\pi_{jl})}{h^{3}}+\frac{\sup_{l\in[L_{n}]}\mbox{Var}\Big{(}\mathbb{I}(\tilde{F}_{j}({\bf x}_{j})\in[b_{l-1},b_{l}))\Big{)}}{h^{3}}+\frac{h^{2}}{2}\cdot({{c_{j}^{*}}^{\prime}}(F^{*}_{j}({\bf x}_{j})|{\bf x}_{-j}))^{2}+\mathcal{O}(h^{3}).$
Then, under Assumption 2 and 3,
$\mbox{MSE}\Big{(}\hat{c}_{j}(\tilde{F}_{j}({\bf x}_{j})|{\bf x}_{-j};\theta)\Big{)}\rightarrow 0$
as $n\rightarrow\infty$, which means that $\hat{c}_{j}(\tilde{F}_{j}({\bf x}_{j})|{\bf x}_{-j};\theta)$ is MSE-consistent estimator for $c_{j}^{*}$.
On the other hand, let $\tilde{h}$ be a smoothing bandwidth of $\tilde{p}_{j}$. For every $\epsilon_{1},\epsilon_{2}>0$, there exists a sufficiently large $M>0$ such that

### v2

@@ Results
We attribute the effectiveness of the feature selection downstream task to our emphasis on estimating ‘conditional’ distributions. In Random Forest (Breiman 2001), each node in a decision tree represents a conditional distribution of a variable conditioned on the splits made by the tree up to that node, and the feature importance is determined by the purity of nodes. Therefore, accurately estimating the conditional distribution can lead to higher performance in preserving the feature importance ranking.
Q2. The missing data mechanism within the parentheses refers to the mechanism applied to the training dataset on which MaCoDE was trained. Despite encountering missing data scenarios such as MAR, MCAR, MNARL, and MNARQ, Table 1 illustrates MaCoDE’s ability to generate high-quality synthetic data regarding machine learning utility while handling incomplete training datasets without significant performance degradation. Even in the presence of missing entries, MaCoDE achieves better metric scores than most baseline models in terms of SMAPE, except for TabDDPM. Additionally, concerning the $F_{1}$ score, MaCoDE either competes competitively or achieves a higher score than other baseline models.
@@
MICE .$010_{\pm.001}.$$845_{\pm.019}.$$\mathbf{.040}_{\pm.002}$
GAIN .$019_{\pm.002}.$$633_{\pm.033}.$$\mathbf{.040}_{\pm.002}$
missMDA .$015_{\pm.001}.$$700_{\pm.022}.$$043_{\pm.002}.$
VAEAC .$\underline{.008}_{\pm.001}$$905_{\pm.016}.$$\mathbf{.040}_{\pm.002}$
MIWAE .$\mathbf{.006}_{\pm.000}$$\underline{.952}_{\pm.012}$$043_{\pm.002}.$
not-MIWAE .$\mathbf{.006}_{\pm.000}$$\mathbf{.949}_{\pm.012}$$\underline{.042}_{\pm.002}$
EGC .$\mathbf{.006}_{\pm.000}$${.996}_{\pm.004}.$$058_{\pm.002}.$
MaCoDE(MAR) .$\mathbf{.006}_{\pm.000}$$963_{\pm.009}.$$051_{\pm.003}.$
Table 2: Q3 under MAR at 0.3 missingness. The means and standard errors of the mean across 5 datasets and 10 repeated experiments are reported. $\downarrow$ denotes lower is better. Coverage close to 0.95 indicates better performance. The best value is bolded, and the second best is underlined.
@@
Q3. The missing data mechanism within the parentheses refers to the mechanism applied to the dataset on which MaCoDE was trained. Table 5 indicates that MaCoDE consistently exhibits competitive performance against all baseline models across metrics assessing multiple imputation performances, including bias, coverage, and confidence interval length. This suggests our proposed approach can support multiple imputations for deriving statistically valid inferences from missing data with the MAR mechanism.
@@
Figure 3: Sensitivity analysis with respect to missingness rate using kings dataset is performed for Q2 under MAR. The analysis focuses on machine learning utility. Results are reported as means and standard errors of the mean from 10 repeated experiments, with error bars representing the standard errors.
@@
Sensitivity analysis. We also conducted a sensitivity analysis by varying the missingness rate of kings dataset. Figure 3 illustrates that, in terms of SMAPE, MaCoDE maintains competitive performance even as the missingness rate increases. Concerning the $F_{1}$ score, MaCoDE outperforms other models at missingness rates of 0.1 and 0.3, but its performance declines beyond a missingness rate of 0.5.
@@ 5 Conclusions and Limitations
This paper introduces an approach to generating synthetic data for mixed-type tabular datasets. Our proposed method integrates histogram-based non-parametric conditional density estimation and the MLM-based approach while bridging the theoretical gap between distributional learning and the consecutive multi-class classification task of MLM. Although our primary goal is to generate synthetic data with high MLu, we empirically demonstrate that we achieve high joint statistical fidelity and MLu simultaneously. Furthermore, empirical experiments validate that our proposed model can generate high-quality synthetic tabular datasets in terms of MLu even when incomplete training datasets are given.
Although MaCoDE demonstrates the ability to perform ‘arbitrary’ conditional density estimation by accommodating various combinations of conditioning sets and target variables, and despite empirical results showing its effectiveness in handling diverse distributions of continuous columns and generating high-quality synthetic data, the model has theoretical limitations. Specifically, it is valid under Lipschitz continuity (Assumption 2). Addressing the limitation of accommodating a broader range of continuous distributions is an important direction for future work.
@@ Appendix A Appendix / supplemental material
Without loss of generality, let $\sigma$ be an identity permutation. We equally space grid points on the interval $[0,1]$ such that bins are represented as $[b_{l-1},b_{l}),l=1,\cdots,L$, where $h$ denotes the bin width such that $Lh=1$. We consider arbitrary $j\in I_{C}$ and ${\bf x}\in\mathbb{R}^{p}$, where ${\bf x}$ is given and fixed. For notational simplicity, let ${\bf x}_{-j}\coloneqq({\bf x}_{1},\cdots,{\bf x}_{j-1})$.
Suppose that $\hat{F}_{j}({\bf x}_{j})\in[b_{l-1},b_{l})$. By the mean value theorem, there exists $u^{*}\in(b_{l-1},b_{l})$ such that
@@
And we have
$\Big{|}c^{*}_{j}(\hat{F}({\bf x}_{j})|{\bf x}_{-j})-\hat{c}_{j}(\hat{F}({\bf x}_{j})|{\bf x}_{-j};\theta)\Big{|}$
$=$ $\left|c^{*}_{j}(\hat{F}({\bf x}_{j})|{\bf x}_{-j})-\frac{\pi_{jl}(g({\bf x};F)\odot{\bf m}^{(j)};\theta)}{h}\right|$
$\leq$ $\left|c^{*}_{j}(\hat{F}({\bf x}_{j})|{\bf x}_{-j})-c^{*}_{j}(u^{*}|{\bf x}_{-j})\right|+\left|c^{*}_{j}(u^{*}|{\bf x}_{-j})-\frac{\pi_{jl}(g({\bf x};F)\odot{\bf m}^{(j)};\theta)}{h}\right|$
$\leq$ $|\hat{F}({\bf x}_{j})-u^{*}|\left|\frac{c^{*}_{j}(\hat{F}({\bf x}_{j})|{\bf x}_{-j})-c^{*}_{j}(u^{*}|{\bf x}_{-j})}{\hat{F}({\bf x}_{j})-u^{*}}\right|+\left|\frac{\pi_{jl}^{*}({\bf x}_{-j})}{h}-\frac{\pi_{jl}(g({\bf x};F)\odot{\bf m}^{(j)};\theta)}{h}\right|$
@@
$\leq$ $hK_{j}+\frac{1}{h}\sum_{l=1}^{L}\left|\pi_{jl}^{*}({\bf x}_{-j})-\pi_{jl}(g({\bf x};F)\odot{\bf m}^{(j)};\theta)\right|.$
where the first equality follows from the definition of $\hat{c}_{j}$.
By the Pinsker’s inequality,
$\frac{1}{2}\sum_{l=1}^{L}\left|\pi_{jl}^{*}({\bf x}_{-j})-\pi_{jl}(g({\bf x};F)\odot{\bf m}^{(j)};\theta)\right|$
$\leq$ $\frac{1}{\sqrt{2}}\left(\sum_{l=1}^{L}\pi_{jl}^{*}({\bf x}_{-j})\log\pi_{jl}^{*}({\bf x}_{-j})-\sum_{l=1}^{L}\pi_{jl}^{*}({\bf x}_{-j})\log\pi_{jl}(g({\bf x};F)\odot{\bf m}^{(j)};\theta)\right)^{1/2}$
$\leq$ $\frac{1}{\sqrt{2}}\left(\sum_{l=1}^{L}\pi_{jl}^{*}({\bf x}_{-j})\log\pi_{jl}^{*}({\bf x}_{-j})-\mathbb{E}_{{\bf y}_{j}|{\bf x}_{-j}}\left[\sum_{l=1}^{L}\mathbb{I}({\bf y}_{j}=l)\log\pi_{jl}(g({\bf x};F)\odot{\bf m}^{(j)};\theta)\right]\right)^{1/2},$
where ${\bf y}_{j}|{\bf x}_{-j}$ is a random variable having a categorical distribution such that $\Pr({\bf y}_{j}=l|{\bf x}_{-j})=\Pr(g({\bf x};\hat{F})_{j}=l|{\bf x}_{-j})=\pi_{jl}^{*}({\bf x}_{-j})$ for all $l\in[L]$.
Then,
$\Big{|}c^{*}_{j}(\hat{F}({\bf x}_{j})|{\bf x}_{-j})-\hat{c}_{j}(\hat{F}({\bf x}_{j})|{\bf x}_{-j};\theta)\Big{|}$ (7)
$\leq$ $hK_{j}+\frac{\sqrt{2}}{h}\left(\sum_{l=1}^{L}\pi_{jl}^{*}({\bf x}_{-j})\log\pi_{jl}^{*}({\bf x}_{-j})-\mathbb{E}_{{\bf y}_{j}|{\bf x}_{-j}}\left[\sum_{l=1}^{L}\mathbb{I}({\bf y}_{j}=l)\log\pi_{jl}(g({\bf x};F)\odot{\bf m}^{(j)};\theta)\right]\right)^{1/2}.$
The total variation distance between $p^{*}_{j}(\cdot|{\bf x}_{-j})$ and $\hat{p}_{j}(\cdot|{\bf x}_{-j};\theta)$ is written as
$\text{TV}\Big{(}p^{*}_{j}(\cdot|{\bf x}_{-j}),\hat{p}_{j}(\cdot|{\bf x}_{-j};\theta)\Big{)}$
$=$ $\frac{1}{2}\int_{\mathbb{R}}\Big{|}p^{*}_{j}({\bf x}_{j}|{\bf x}_{-j})-\hat{p}_{j}({\bf x}_{j}|{\bf x}_{-j};\theta)\Big{|}d{\bf x}_{j}$